OpenStudy (anonymous):

HELP Write the equation of the ellipse in standard form. 9x^2+4y^2-72x+40y+208=0

OpenStudy (anonymous):

Rewrite as: 9(x^2 - 8x) +4(y^2 +10y) +208 = 0 Complete squares: 9(x^2 -8x + 16) + 4(y^2 + 10y + 25) + 208 = 244 9(x-4)^2 +4(y+5)^2 = 36 Divide by 36: (x-4)^2/4 + (y+5)^2/9 = 1 x^2/9 + y^2/36 = 1 Center at(0,0) Major axis || to y-axis y-intercepts at +/- 6 x-intercepts at +/- 3

OpenStudy (anonymous):

was i correct

OpenStudy (anonymous):

i was...

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

@72595210002

OpenStudy (anonymous):

D

OpenStudy (anonymous):

does it elp

OpenStudy (anonymous):

thank you honey!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

can u help me